OpenStudy (anonymous):

123x^2 - 492x +813 = 444 what is x

OpenStudy (solomonzelman):

first subtract 813 from both sides.

OpenStudy (anonymous):

ok i did that

OpenStudy (amoodarya):

if you can simplify first simplify then will have two roots x=1,3

OpenStudy (solomonzelman):

Divide everything by 123.

OpenStudy (solomonzelman):

see? they just made it look confusing, but you can take 123 out of all this.

OpenStudy (solomonzelman):

123x^2+492x=-369 divide everything by 123.

OpenStudy (solomonzelman):

x^2+4x=-3 you can add 3 to both sides to get x^2+4x+3=0 factor (x+3)(x+1)=0 x=-3 or x=-1